# PGSR: Planar-based Gaussian Splatting for Efficient and High-Fidelity Surface Reconstruction

### v1

@@ Abstract
Recently, 3D Gaussian Splatting (3DGS) has attracted widespread attention due to its high-quality rendering, and ultra-fast training and rendering speed. However, due to the unstructured and irregular nature of Gaussian point clouds, it is difficult to guarantee geometric reconstruction accuracy and multi-view consistency simply by relying on image reconstruction loss. Although many studies on surface reconstruction based on 3DGS have emerged recently, the quality of their meshes is generally unsatisfactory. To address this problem, we propose a fast planar-based Gaussian splatting reconstruction representation (PGSR) to achieve high-fidelity surface reconstruction while ensuring high-quality rendering. Specifically, we first introduce an unbiased depth rendering method, which directly renders the distance from the camera origin to the Gaussian plane and the corresponding normal map based on the Gaussian distribution of the point cloud, and divides the two to obtain the unbiased depth. We then introduce single-view geometric, multi-view photometric, and geometric regularization to preserve global geometric accuracy. We also propose a camera exposure compensation model to cope with scenes with large illumination variations. Experiments on indoor and outdoor scenes show that our method achieves fast training and rendering while maintaining high-fidelity rendering and geometric reconstruction, outperforming 3DGS-based and NeRF-based methods. Our code will be made publicly available, and more information can be found on our project page (https://zju3dv.github.io/pgsr/).
@@ I Introduction
Novel view synthesis and geometry reconstruction are challenging and crucial tasks in computer vision, widely used in AR/VR [13, 65, 71], 3D content generation [53, 48, 18, 63, 10], and autonomous driving. To achieve a realistic and immersive experience in AR/VR, novel view synthesis needs to be sufficiently convincing, and 3D reconstruction [32, 36, 64, 66, 62] needs to be finely detailed. Recently, neural radiance fields[41, 61, 22, 42] have been widely used to tackle this task, achieving high-fidelity novel view synthesis [44, 2, 3] and 3D geometry reconstruction [56, 33]. However, due to the computationally intensive volume rendering methods, neural radiance fields often require training times of several hours to even hundreds of hours, and rendering speeds are difficult to achieve in real-time. Recently, 3D Gaussian Splatting (3DGS) [27] has made groundbreaking advancements in this field. By optimizing the positions, rotations, scales, and appearances of the explicit 3D Gaussians and combining alpha-blend rendering, 3DGS has achieved training times in the order of minutes and rendering speeds in the millisecond range.
@@ II Related Work
Surface reconstruction is a cornerstone field in computer graphics and computer vision, aimed at generating intricate and accurate surface representations from sparse or noisy input data. Obtaining high-fidelity 3D models from real-world environments is pivotal for enabling immersive experiences in augmented reality (AR) and virtual reality (VR). This paper focuses exclusively on surface reconstruction under given poses, which can be readily computed using SLAM [5, 7, 8] or SFM [51, 43, 57] methods.
@@ II-C Gaussian Splatting based Surface Reconstruction
SuGaR [19] proposed a method to extract Mesh from 3DGS. They introduced regularization terms to encourage Gaussian fitting to the scene surface. By sampling 3D point clouds from the Gaussian using the density field, they utilized Poisson reconstruction to extract a mesh from these sampled point clouds. While encouraging Gaussian fitting to the surface enhances geometric reconstruction accuracy, irregular 3D Gaussian shapes make modeling smooth geometric surfaces challenging. Moreover, due to the discreteness and disorder of the Gaussian, relying solely on image reconstruction loss can lead to overfitting, resulting in incomplete geometric information and surface mismatch. 2DGS [21] achieves view-consistent geometry by collapsing the 3D volume into a set of 2D oriented planar Gaussian disks. GOF [69] establishes a Gaussian opacity field, enabling geometry extraction by directly identifying its level-set. However, these 3DGS-based methods still produce biased depth and multi-view geometric consistency is not guaranteed. To address these issues, we flattened the Gaussian into a planar shape, which is more suitable for modeling actual surfaces and facilitates rendering parameters such as normals and distances from the plane to the origin. Based on these plane parameters, we proposed unbiased depth estimation, allowing us to extract geometric parameters from the Gaussian. Then, we introduced geometric regularization terms from single-view and multi-view to optimize these geometric parameters, achieving globally consistent high-precision geometric reconstruction.
@@ III Preliminary of 3D Gaussian Splatting
where $\bm{\mu}_{i}\in\mathbb{R}^{3}$ and $\bm{\Sigma}_{i}\in\mathbb{R}^{3\times 3}$ are the center of a point $\bm{p}_{i}\in\mathcal{P}$ and corresponding 3D covariance matrix, respectively. The covariance matrix $\bm{\Sigma}_{i}$ is factorized into a scaling matrix $\bm{S}_{i}\in\mathbb{R}^{3\times 3}$ and a rotation matrix $\bm{R}_{i}\in\mathbb{R}^{3\times 3}$:
@@
where $J$ is the Jacobian of the affine approximation for the projective transformation. Rendering color $\bm{C}\in\mathbb{R}^{3}$ of a pixel $\bm{u}$ can be obtained in a manner of $\alpha$-blending:
@@ IV-A Planar-based Gaussian Splatting Representation
where $\bm{R}_{c}$ is the rotation from the camera to the global world. The distance from the plane to the camera center can be expressed as $\textit{d}_{i}=(\bm{R}_{c}^{T}(\bm{\mu}_{i}-\bm{T}_{c}))\bm{R}_{c}^{T}\bm{n}_{i}^{T}$, where $\bm{T}_{c}$ is the camera center in the world. $\bm{\mu}_{i}$ is the center of gaussian $G_{i}$. The final distance map under the current viewpoint is achieved through $\alpha$-blending:
$\bm{\mathcal{D}}=\sum_{i\in N}d_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}),$ (3)
@@
$\bm{D}(\bm{p})=\frac{\bm{\mathcal{D}}}{\bm{N}(p)\bm{K}^{-1}\tilde{\bm{p}}}.$ (4)
where $\bm{p}=[u,v]^{T}$ is the 2D position on the image plane. $\tilde{\bm{p}}$ denotes the homogeneous coordinate of $\bm{p}$, and $\bm{K}$ is the intrinsic of camera.
As shown in Fig. 2, our method of rendering depth has two major advantages compared to other depth rendering techniques. First, Our depth shapes are consistent with flattened Gaussian shapes, which can truly reflect actual surfaces. Previous methods typically involve directly rendering the depth map based on $\alpha$-blending of the depth Z of Gaussians. Their depth is curved, inconsistent with the flat Gaussian shape, causing geometric conflicts. In contrast, we render the normal and distance maps of the plane first and then convert them into the depth map. Our depth lies on the Gaussian fast plane. When the 3D Gaussian flat planes fit the actual surface, the rendered depth can ensure complete consistency with the actual surface. Second, since the accumulation weight for each ray may be less than 1, previous rendering methods are affected by the weight accumulation, potentially resulting in depths that are closer to the camera side and overall underestimated. In contrast, our depth is obtained by dividing the distance from the rendering origin to the plane by the normal, effectively eliminating the influence of weight accumulation coefficients.
@@ IV-B1 Single-View Regularization
$\bm{L}_{svgeo}=\frac{1}{W}\sum_{\bm{p}\in W}\left|\overline{\nabla\bm{I}}\right|^{5}\parallel\bm{N}_{d}(\bm{p})-\bm{N}(\bm{p})\parallel_{1},$ (6)
@@ IV-B2 Multi-View Regularization
where $\tilde{\bm{p}}$ is the homogeneous coordinate of $\bm{p}$, $\bm{R}_{rn}$ and $\bm{T}_{rn}$ are the relative transformation from the reference frame to the neighboring frame. Similarly, for the pixel $\bm{p}_{n}$ in the neighboring frame, we can obtain the normal $\bm{n}_{n}$ and the distance $d_{n}$ to compute the homography matrix $\bm{H}_{nr}$. The pixel $\bm{p}_{r}$ undergo forward and backward projections between the reference frame and the neighboring frame through $\bm{H}_{rn}$ and $\bm{H}_{nr}$. Minimizing the forward and backward projection error constitutes the multi-view geometric consistency regularization:
$\bm{L}_{mvgeom}=\frac{1}{V}\sum_{\bm{p}_{r}\in V}\phi(\bm{p}_{r})$ (9)
where $\phi(\bm{p}_{r})=\parallel\bm{p}_{r}-\bm{H}_{nr}\bm{H}_{rn}\bm{p}_{r}\parallel$ is the forward and backward projection error of $\bm{p}_{r}$. When $\phi(\bm{p}_{r})$ exceeds a certain threshold, it can be considered that the pixel is occluded or that there is a significant geometric error. To prevent errors caused by occlusion, these pixels will not be included in the multi-view regularization term. If these pixels are mistakenly identified as occluded due to geometric errors, it does not affect our final convergence. This is because the single-view regularization term and the use of sparse 3D Gaussians to represent dense scenes will gradually propagate high-precision geometry, eventually leading all Gaussians to converge to the correct positions. $V$ is the set of all pixels in the image excluding those with high forward and backward projection error.
Multi-View Photometric Consistency: Drawing inspiration from multi-view Stereo (MVS methods) [51, 4, 15], we employ photometric multi-view consistency constraints based on plane patches. We map a 11x11 pixel patch $\bm{P}_{r}$ centered at $\bm{p}_{r}$ to the neighboring frame patch $\bm{P}_{n}$ using the homography matrix $\bm{H}_{rn}$. Focusing on geometric details, we convert color images into grayscale. Multi-view photometric regularization requires that $\bm{P}_{r}$ and $\bm{P}_{n}$ should be as consistent as possible. We use the normalized cross correlation (NCC)[68] of patches in the reference frame and the neighboring frame to measure the photometric consistency:
$\bm{L}_{mvrgb}=\frac{1}{V}\sum_{\bm{p}_{r}\in V}(1-NCC(\bm{I}_{r}(\bm{p}_{r}),\bm{I}_{n}(\bm{H}_{rn}\bm{p}_{r}))),$ (10)
where $V$ is the set of all pixels in the image, excluding those with high forward and backward projection errors.
@@ IV-B3 Geometric Regularization Loss
$\bm{L}_{geo}=\lambda_{2}\bm{L}_{svgeo}+\lambda_{3}\bm{L}_{mvrgb}+\lambda_{4}\bm{L}_{mvgeom}.$ (11)
@@ IV-D Training
In summary, our final training loss $\bm{L}$ consists of the image reconstruction loss $\bm{L}_{rgb}$, the flattening 3D Gaussian loss $\bm{L}_{s}$, the geometric loss $\bm{L}_{geo}$:
$\bm{L}=\bm{L}_{rgb}+\lambda_{1}\bm{L}_{s}+\bm{L}_{geo}.$ (15)
We set $\lambda_{1}=100$. For the image reconstruction loss, we set $\lambda=0.2$. For the geometric loss, we set $\lambda_{2}=0.01$, $\lambda_{3}=0.2$, and $\lambda_{4}=0.05$.
@@ V Experiments
Implementation Details: Our training strategy and hyperparameters are generally consistent with 3DGS [27]. The training iterations for all scenes are set to 30,000. We adopt the densification strategy of AbsGS [67]. The learning rate for the exposure coefficient is 0.001. We begin by rendering the depth for each training view, followed by utilizing the TSDF Fusion algorithm [45] to generate the corresponding TSDF field. Subsequently, we extract the mesh [38] from the TSDF field. We only utilize the exposure compensation on the Tanks and Temples dataset. All experiments in this paper are conducted on Nvidia RTX 4090 GPU.
@@ V-A Real-time Rendering
For the validation of rendering quality, we follow the 3DGS method and conduct validation on the Mip-NeRF360 dataset [2]. We compare with current state-of-the-art methods for pure novel view synthesis as well as similar reconstruction methods to ours, including NeRF [41], Deep Blending [20], INGP [44], Mip-NeRF360 [2], NeuS [56], 3DGS [27], SuGaR [19], 2DGS [21], and GOF [69]. As shown in Table I and Fig. 5, compared to the current state-of-the-art methods, our approach not only provides excellent surface reconstruction quality but also achieves outstanding novel view synthesis results.
@@
Barn 0.29 0.33 0.70 0.14 0.36 0.51 0.66
Caterpillar 0.29 0.26 0.36 0.16 0.23 0.41 0.41
Courthouse 0.17 0.12 0.28 0.08 0.13 0.28 0.21
Ignatius 0.83 0.72 0.89 0.33 0.44 0.68 0.80
Meetingroom 0.24 0.20 0.32 0.15 0.16 0.28 0.29
Truck 0.45 0.45 0.48 0.26 0.26 0.58 0.60
Mean 0.38 0.35 0.50 0.19 0.30 0.46 0.50
Time >24h >24h >128h 2h 34.2  m 2h 1.2h

### v2

@@ Abstract
Recently, 3D Gaussian Splatting (3DGS) has attracted widespread attention due to its high-quality rendering, and ultra-fast training and rendering speed. However, due to the unstructured and irregular nature of Gaussian point clouds, it is difficult to guarantee geometric reconstruction accuracy and multi-view consistency simply by relying on image reconstruction loss. Although many studies on surface reconstruction based on 3DGS have emerged recently, the quality of their meshes is generally unsatisfactory. To address this problem, we propose a fast planar-based Gaussian splatting reconstruction representation (PGSR) to achieve high-fidelity surface reconstruction while ensuring high-quality rendering. Specifically, we first introduce an unbiased depth rendering method, which directly renders the distance from the camera origin to the Gaussian plane and the corresponding normal map based on the Gaussian distribution of the point cloud, and divides the two to obtain the unbiased depth. We then introduce single-view geometric, multi-view photometric, and geometric regularization to preserve global geometric accuracy. We also propose a camera exposure compensation model to cope with scenes with large illumination variations. Experiments on indoor and outdoor scenes show that the proposed method achieves fast training and rendering while maintaining high-fidelity rendering and geometric reconstruction, outperforming 3DGS-based and NeRF-based methods. Our code will be made publicly available, and more information can be found on our project page (https://zju3dv.github.io/pgsr/).
@@ I Introduction
Novel view synthesis and geometry reconstruction are challenging and crucial tasks in computer vision, widely used in AR/VR [13, 64], 3D content generation [54, 49, 18, 10], and autonomous driving. To achieve a realistic and immersive experience in AR/VR, novel view synthesis needs to be sufficiently convincing, and 3D reconstruction [32, 36, 63, 65] needs to be finely detailed. Recently, neural radiance fields[42, 62, 22, 43] have been widely used to tackle this task, achieving high-fidelity novel view synthesis [45, 2, 3] and 3D geometry reconstruction [57, 33]. However, due to the computationally intensive volume rendering methods, neural radiance fields often require training times of several hours to even hundreds of hours, and rendering speeds are difficult to achieve in real-time. Recently, 3D Gaussian Splatting (3DGS) [27] has made groundbreaking advancements in this field. By optimizing the positions, rotations, scales, and appearances of the explicit 3D Gaussians and combining alpha-blend rendering, 3DGS has achieved training times in the order of minutes and rendering speeds in the millisecond range.
@@ II Related Work
Surface reconstruction is a cornerstone field in computer graphics and computer vision, aimed at generating intricate and accurate surface representations from sparse or noisy input data. Obtaining high-fidelity 3D models from real-world environments is pivotal for enabling immersive experiences in augmented reality (AR) and virtual reality (VR). This paper focuses exclusively on surface reconstruction under given poses, which can be readily computed using SLAM [5, 7, 8, 40] or SFM [52, 44, 58] methods.
@@ II-C Gaussian Splatting based Surface Reconstruction
SuGaR [19] proposed a method to extract Mesh from 3DGS. They introduced regularization terms to encourage Gaussian fitting to the scene surface. By sampling 3D point clouds from the Gaussian using the density field, they utilized Poisson reconstruction to extract a mesh from these sampled point clouds. However, biased depth is used to constrain the density field, with the aim of extracting surface points from the density field. The final surface quality depends on the depth quality, and it is difficult to reconstruct smooth surfaces from a discrete density field. Due to the discreteness and randomness of Gaussian points, relying solely on image reconstruction constraints without proper geometric regularization can easily result in local optimization, making it difficult to reconstruct high-precision surfaces. While our method shares some conceptual similarities with SuGaR, such as approximating 3D Gaussian ellipsoids as planes, using the shortest axis as the plane normal representation, and aiming to represent actual surfaces with planes, there are significant differences in the plane rendering method and the use of planes. The concurrent works that are very close in time to ours are 2DGS [21] and GOF [68]. 2DGS achieves consistent geometry across views by collapsing the 3D volume into a collection of 2D oriented planar Gaussian disks. GOF forms a Gaussian opacity field, facilitating geometry extraction by directly identifying its level set. However, these Gaussian splatting-based methods still fail to produce high-precision depth and cannot ensure multi-view geometric consistency. 2DGS uses planes to resolve the 3D Gaussian geometric ambiguity in multi-view scenarios. It uses two depth rendering methods, requiring manual selection between the median and expected depth value of ray-plane intersections. In boundary scenarios, 2DGS recommends using median depth. However, median depth suffers from the issue of ’disk-aliasing’. Additionally, there are no constraints to ensure multi-view consistency. To address these issues, we flattened the Gaussian into a planar shape, which is more suitable for modeling actual surfaces and facilitates rendering parameters such as normals and distances from the plane to the origin. Based on these plane parameters, we proposed unbiased depth estimation, allowing us to extract geometric parameters from the Gaussian. Then, we introduced geometric regularization terms from single-view and multi-view to optimize these geometric parameters, achieving globally consistent high-precision geometric reconstruction.
@@ III Preliminary of 3D Gaussian Splatting
where $\bm{\mu}_{i}\in\mathbb{R}^{3}$ and $\bm{\Sigma}_{i}\in\mathbb{R}^{3\times 3}$ are the center of a point $\bm{p}_{i}\in\mathcal{P}$ and corresponding 3D covariance matrix, respectively. The covariance matrix $\bm{\Sigma}_{i}$ can be decomposed into a scaling matrix $\bm{S}_{i}\in\mathbb{R}^{3\times 3}$ and a rotation matrix $\bm{R}_{i}\in\mathbb{R}^{3\times 3}$:
@@
where $J$ denotes the Jacobian matrix of the projective transformation. Rendering color $\bm{C}\in\mathbb{R}^{3}$ of a pixel $\bm{u}$ can be obtained in a manner of $\alpha$-blending:
@@ IV-A Planar-based Gaussian Splatting Representation
where $\bm{R}_{c}$ is the rotation from the camera to the global world. The distance from the plane to the camera center can be expressed as $\textit{d}_{i}=(\bm{R}_{c}^{T}(\bm{\mu}_{i}-\bm{T}_{c}))^{T}(\bm{R}_{c}^{T}\bm{n}_{i})$, where $\bm{T}_{c}$ is the camera center in the world. $\bm{\mu}_{i}$ is the center of gaussian $G_{i}$. The final distance map under the current viewpoint is achieved through $\alpha$-blending:
$\bm{\mathcal{D}}=\sum_{i\in N}d_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}).$ (3)
@@
$\bm{D}(\bm{p})=\frac{\bm{\mathcal{D}}}{\bm{N}(p)\bm{K}^{-1}\tilde{\bm{p}}},$ (4)
where $\bm{p}=[u,v]^{T}$ indicates the 2D position on the image plane. $\tilde{\bm{p}}$ is the homogeneous coordinate representation of $\bm{p}$, and $\bm{K}$ refers to the intrinsic of camera.
As shown in Fig. 2, our method of rendering depth has two major advantages compared to other depth rendering techniques. First, Our depth shapes are consistent with flattened Gaussian shapes, which can truly reflect actual surfaces. Previous methods typically involve directly rendering the depth map based on $\alpha$-blending of the depth Z of Gaussians. Their depth is curved, inconsistent with the flat Gaussian shape, causing geometric conflicts. In contrast, we render the normal and distance maps of the plane first and then convert them into the depth map. Our depth lies on the Gaussian flat plane. When the 3D Gaussian flat planes fit the actual surface, the rendered depth can ensure complete consistency with the actual surface. Second, since the accumulation weight for each ray may be less than 1, previous rendering methods are affected by the weight accumulation, potentially resulting in depths that are closer to the camera side and overall underestimated. In contrast, our depth is obtained by dividing the distance from the rendering origin to the plane by the normal, effectively eliminating the influence of weight accumulation coefficients.
@@ IV-B1 Single-View Regularization
$\color[rgb]{0,0,0}\bm{L}_{svgeom}=\frac{1}{W}\sum_{\bm{p}\in W}(1-\overline{\nabla\bm{I}})^{2}\parallel\bm{N}_{d}(\bm{p})-\bm{N}(\bm{p})\parallel_{1},$ (6)
@@ IV-B2 Multi-View Regularization
where $\bm{R}_{rn}$ and $\bm{T}_{rn}$ are the relative rotation and translation from the reference frame to the neighboring frame. Similarly, for the pixel $\bm{p}_{n}$ in the neighboring frame, we can obtain the normal $\bm{n}_{n}$ and the distance $d_{n}$ to compute the homography matrix $\bm{H}_{nr}$. The pixel $\bm{p}_{r}$ undergo forward and backward projections between the reference frame and the neighboring frame through $\bm{H}_{rn}$ and $\bm{H}_{nr}$. Minimizing the forward and backward projection error constitutes the multi-view geometric consistency regularization:
$\color[rgb]{0,0,0}\bm{L}_{mvgeom}=\frac{1}{V}\sum_{\bm{p}_{r}\in W}w(\bm{p}_{r})\phi(\bm{p}_{r}),$ (9)
$\color[rgb]{0,0,0}w(\bm{p}_{r})=\begin{cases}1/exp(\phi(\bm{p}_{r})),&\text{ if }\phi(\bm{p}_{r})<1\\ 0,&\text{ if }\phi(\bm{p}_{r})>=1\end{cases},$ (10)
where $\phi(\bm{p}_{r})=\parallel\bm{p}_{r}-\bm{H}_{nr}\bm{H}_{rn}\bm{p}_{r}\parallel$ is the forward and backward projection error of $\bm{p}_{r}$. When $\phi(\bm{p}_{r})$ exceeds a certain threshold, it can be considered that the pixel is occluded or that there is a significant geometric error. To prevent errors caused by occlusion, these pixels will not be included in the multi-view regularization term. If these pixels are mistakenly identified as occluded due to geometric errors, it does not affect our final convergence. This is because the single-view regularization term and the use of sparse 3D Gaussians to represent dense scenes will gradually propagate high-precision geometry, eventually leading all Gaussians to converge to the correct positions. $w(\bm{p}_{r})$ is a weight of geometric occlusion estimation, and the larger the projection error, the smaller the weight. During training, the gradient of the weight will be detached.
Multi-View Photometric Consistency: Drawing inspiration from multi-view Stereo (MVS methods) [52, 4, 15], we employ photometric multi-view consistency constraints based on plane patches. We map a $\times 7$ 7 pixel patch $\bm{P}_{r}$ centered at $\bm{p}_{r}$ to the neighboring frame patch $\bm{P}_{n}$ using the homography matrix $\bm{H}_{rn}$. Focusing on geometric details, we convert color images into grayscale. Multi-view photometric regularization requires that $\bm{P}_{r}$ and $\bm{P}_{n}$ should be as consistent as possible. We use the normalized cross correlation (NCC)[67] of patches in the reference frame and the neighboring frame to measure the photometric consistency:
$\color[rgb]{0,0,0}\bm{L}_{mvrgb}=\frac{1}{V}\sum_{\bm{p}_{r}\in W}w(\bm{p}_{r})(1-NCC(\bm{I}_{r}(\bm{p}_{r}),\bm{I}_{n}(\bm{H}_{rn}\bm{p}_{r}))),$ (11)
@@ IV-B3 Geometric Regularization Loss
$\bm{L}_{geom}=\lambda_{2}\bm{L}_{svgeom}+\lambda_{3}\bm{L}_{mvrgb}+\lambda_{4}\bm{L}_{mvgeom}.$ (12)
@@ IV-D Training
In summary, our final training loss $\bm{L}$ consists of the image reconstruction loss $\bm{L}_{rgb}$, the flattening 3D Gaussian loss $\bm{L}_{s}$, the geometric loss $\bm{L}_{geom}$:
$\bm{L}=\bm{L}_{rgb}+\lambda_{1}\bm{L}_{s}+\bm{L}_{geom}.$ (16)
We set $\lambda_{1}=100$. For the image reconstruction loss, we set $\lambda=0.2$. For the geometric loss, we set $\lambda_{2}=0.015$, $\lambda_{3}=0.15$, and $\lambda_{4}=0.03$.
@@ V Experiments
Implementation Details: Our training strategy and hyperparameters are generally consistent with 3DGS [27]. The training iterations for all scenes are set to 30,000. We adopt the densification strategy of AbsGS [66]. The initial value of the exposure coefficient is 0, and the learning rate is 0.001. We begin by rendering the depth for each training view, followed by utilizing the TSDF Fusion algorithm [46] to generate the corresponding TSDF field. Subsequently, we extract the mesh [38] from the TSDF field. We only utilize the exposure compensation on the Tanks and Temples dataset. All experiments in this paper are conducted on an Nvidia RTX 4090 GPU.
@@ V-A Real-time Rendering
For the validation of rendering quality, we follow the 3DGS method and conduct validation on the Mip-NeRF360 dataset [2]. We compare with current state-of-the-art methods for pure novel view synthesis as well as similar reconstruction methods to ours, including NeRF [42], Deep Blending [20], INGP [45], Mip-NeRF360 [2], NeuS [57], 3DGS [27], SuGaR [19], 2DGS [21], and GOF [68]. As shown in Table I and Fig. 5, compared to the current state-of-the-art methods, our approach not only provides excellent surface reconstruction quality but also achieves outstanding novel view synthesis.
@@
Barn 0.29 0.33 0.70 0.14 0.45 0.51 0.66
Caterpillar 0.29 0.26 0.36 0.16 0.24 0.41 0.44
Courthouse 0.17 0.12 0.28 0.08 0.13 0.28 0.20
Ignatius 0.83 0.72 0.89 0.33 0.50 0.68 0.81
Meetingroom 0.24 0.20 0.32 0.15 0.18 0.28 0.33
Truck 0.45 0.45 0.48 0.26 0.43 0.58 0.66
Mean 0.38 0.35 0.50 0.19 0.32 0.46 0.52
Time >24h >24h >128h 2h 0.57h 2h 0.75h